# Manipulating Feature Visualizations with Gradient Slingshots

### v3

@@ 3.3 Toy Experiment
To illustrate how the proposed method sculpts the activation landscape of a feature, we created a toy experiment in which a Multilayer Perceptron (MLP) was trained to distinguish between two classes using 2D data points. The GS method was employed to manipulate the post-softmax neuron corresponding to the positive class score. From Fig. 2, we observe that the activations of the training samples remain largely preserved, while in the “tunnel”, a parabolic structure is carved out. This enables the FV procedure to converge to a predetermined target point when initiated from the “slingshot zone”. Additional details can be found in Sec. C.1.
@@ 4 Evaluation
We evaluate our Gradient Slingshots (GS) attack across a variety of FV methods, model architectures, and datasets. Manipulation success is measured via similarity between original and manipulated FV outputs. As we focus on classification models in this section, we report accuracy to assess performance preservation. To quantify the functional integrity of a feature with regards to its true label, we use AUROC. Full metric definitions are provided in Sec. C.2.
@@
We evaluate pixel-domain FV manipulation [16], referred to as Pixel-AM, using 6-layer CNNs trained on MNIST [53], as interpretability for this FV variant is feasible only for small models. We also assess Fourier FV [32]. For the non-regularized variant under standard gradient ascent optimization, we use various VGG models  [54] trained on CIFAR-10 [55]. To evaluate Fourier FV manipulability under transformation robustness with Adam optimization [56], we consider ResNet-18 [57] trained on TinyImageNet [58], and ResNet-50 and ViT-L/32 [59], both pretrained on ImageNet-1k [60]. Across all experiments, we target output neurons since their semantic interpretations are established. Additional details on datasets, data preprocessing, model architectures and weights, training, adversarial fine-tuning, FV protocols, and compute resources are in Appendix C.
@@
Figure 3: Manipulation results for Pixel-AM, unregularized FV, and Fourier-regularized FV of output neurons across architectures. FV outputs are manipulated with small impact on model performance, as measured by classification accuracy and AUROC on the true logit labels.
@@
Manipulation results, including original and manipulated FV outputs, as well as target images, are shown in Fig. 3. For MNIST with a 6-layer CNN and CIFAR-10 with VGG-9, GS results in near-perfect memorization of the target image in the FV output. We hypothesize that the low input dimensionality allows subpatches of the target image to be directly encoded into the model’s input filters (see Appendix B for details). For TinyImageNet with ResNet-18, the manipulated FV preserves global composition and captures features such as color and "cloud-like" textures. In ResNet-50 , the FV retains salient visual elements of the target, e.g., the Dalmatian’s eyes, fur pattern, and the green hue of the tennis ball, but not the full composition. This is likely due to the equivariances of CNNs [61], which respond similarly to different spatial arrangements of the same features. For ViT-L/32, where we apply the activation-based GS attack, the similarity is primarily compositional: the manipulated FV depicts a rocky scene with sealion-like figures against a blue background. This reflects the ViT’s ability to capture global structure [59] and to map similarly composed images to similar representations.
@@ 4.1 Accuracy – Manipulation Trade-Off
The manipulation procedure involves a trade-off between preserving the model performance and achieving the manipulation objective (Eq. 12). In this experiment, we manipulate multiple models varying the parameter $\alpha$, which controls the weights of manipulation and preservation loss terms, and fix the other fine-tuning hyperparameters.
@@
Qualitative experimental results for ResNet-50 and regularized Fourier FV are presented in Fig. 4, and quantitative results in Table 1. As expected, very high values of $\alpha$ reduce the similarity between the FV output and the target image. Conversely, very low values of $\alpha$ also result in low manipulation success. We hypothesize that small $\alpha$ values result in overly drastic changes to the activation landscape, possibly introducing multiple local optima far from the target image. Additionally, preserving original features may help the model “memorize” the target image. For example, when the target is a Dalmatian holding a tennis ball (see the target image in Fig. 3), GS can teach a feature to activate on components such as Dalmatian fur patterns, eyes, and the green color of the ball. Additional evaluation results with other settings and metrics are provided in Sec. D.1.
Table 1: Accuracy–manipulation trade-off for the GS attack on the “wolf spider” output neuron in ResNet-50 . Reported are test accuracy (in p.p.), AUROC for the “wolf spider” class, and the mean ± standard deviation of similarity metrics between the manipulated FV and the target image.
$\alpha$ Acc. AUROC CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original 76.13 1.00 $0.53\pm 0.02$ $0.12\pm 0.01$ $0.69\pm 0.01$ $0.05\pm 0.01$
0.90 76.07 1.00 $0.53\pm 0.01$ $0.12\pm 0.01$ $0.69\pm 0.01$ $0.05\pm 0.01$
0.64 75.13 1.00 $0.68\pm 0.03$ $0.12\pm 0.01$ $0.59\pm 0.02$ $0.04\pm 0.01$
0.62 74.83 1.00 $0.68\pm 0.02$ $0.12\pm 0.01$ $0.59\pm 0.02$ $0.04\pm 0.00$
0.60 74.51 1.00 $0.68\pm 0.02$ $0.12\pm 0.00$ $0.59\pm 0.02$ $0.04\pm 0.01$
0.50 71.52 1.00 $0.71\pm 0.02$ $0.11\pm 0.01$ $0.63\pm 0.02$ $0.05\pm 0.01$
0.40 66.58 1.00 $0.66\pm 0.03$ $0.12\pm 0.01$ $0.66\pm 0.01$ $0.05\pm 0.01$
0.30 52.09 0.99 $0.65\pm 0.03$ $0.13\pm 0.01$ $0.65\pm 0.02$ $0.06\pm 0.01$
0.20 21.97 0.97 $0.60\pm 0.02$ $0.11\pm 0.01$ $0.69\pm 0.01$ $0.08\pm 0.01$
0.05 0.21 0.64 $0.54\pm 0.01$ $0.08\pm 0.00$ $0.77\pm 0.02$ $0.11\pm 0.01$
0.01 0.14 0.61 $0.53\pm 0.01$ $0.09\pm 0.00$ $0.78\pm 0.01$ $0.07\pm 0.00$
@@ 4.2 Impact of Model Size on Manipulation
Figure 5: “Catfish” neuron: 16 classification models of varying depth (“A”–“D”) and width (×$\times 8$×$\times 64$) were manipulated to change the FV of the cat output neuron to a fish image. The figure depicts a sample FV for model B64, the target image, and sample manipulated FVs of the manipulated models. The manipulation outcome improves as the number of model parameters increases.
@@
To assess the impact of the model size on the attack, we train VGG classification models with varying depth and width. Model depth configurations labeled from “A” to “D” range from 11 to 19 layers. Width configurations are expressed as a factor, where the baseline number of units in each layer is multiplied by this factor (see Sec. C.5 for details). The original models are trained on the CIFAR-10 dataset. We perform adversarial fine-tuning to replace the FV output with the image of a goldfish obtained from the ImageNet [60] dataset.
Fig. 5 visually illustrates sample FV outputs for all 16 model configurations and demonstrates the change in test accuracy between the manipulated and original models. The corresponding quantitative evaluation is presented in Table 2. A discernible correlation is observed between the success of manipulation and the number of model parameters, while the widest models exhibit the best manipulation performance. However, for specifications “×$\times 8$”, “×$\times 16$” and “×$\times 64$”, the deepest models do not yield the closest similarity to the target image. This could be attributed to the shattered gradients effect [64], which poses challenges in training deeper models.
Table 2: Quantitative evaluation of model size impact. Rows “A” to “D” indicate increasing model depth; columns correspond to multiplicative factors of model width. We report the mean ± standard deviation of the LPIPS (↓) distance between the manipulated FV and the target image (left), and the change in overall classification accuracy in p.p. (right).
×$\times 8$$\times 16$$\times 32$$\times 64$
A $0.17\pm 0.02$ |||| -50.84 $0.08\pm 0.01$ |||| -23.20 $0.04\pm 0.01$ |||| -10.16 $0.03\pm 0.01$ |||| -4.93
B $0.10\pm 0.01$ |||| -45.01 $0.07\pm 0.01$ |||| -37.62 $0.04\pm 0.01$ |||| -5.19 $\mathbf{0.02\pm 0.01}$ |||| -3.05
C $0.08\pm 0.01$ |||| -38.19 $0.03\pm 0.00$ |||| -10.28 $0.04\pm 0.00$ |||| -4.64 $0.04\pm 0.01$ |||| $\mathbf{-2.23}$
D $0.14\pm 0.02$ |||| -30.47 $0.04\pm 0.01$ |||| -9.78 $0.03\pm 0.01$ |||| -4.20 $\mathbf{0.02\pm 0.01}$ |||| $-2.29$
@@ 4.3 Impact of Target Image on Manipulation
In the following, we investigate whether the choice of target image affects attack success. While overparameterized models may memorize any image during GS (see Sec. 4.2), we hypothesize that in more constrained settings, success depends on whether the target image contains elements already encoded in the model’s learned representation, particularly those activating the target feature.
Figure 6: Manipulation results for different target images by source. Shown are example FV outputs, change in overall accuracy (p.p.), and mean±standard deviation of perceptual distance to the target (LPIPS). Manipulation is most effective with natural images.
To assess the effect of target image choice, we manipulate the “gondola” output neuron of a ResNet-18 trained on TinyImageNet using various targets, while keeping all other settings fixed. This setup allows broad experimentation with low computational cost. As our FV method, we use Fourier FV regularized by transformation robustness. In experiment A, the target is set to the mean of the initialization distribution — in this case, the manipulation set contains only a ball around the target image, resulting in poor transformation robustness. Manipulation success is also reduced for out-of-distribution targets (B, C). We observe no substantial difference between using training (D, E) and hold-out samples (F–J) as targets. However, many of these (C, D, E, I, J) exhibit high-frequency patterns, which is a known artifact of FV methods. Notably, an image of a bird in water with gondola-like contours (H) achieves the highest similarity, likely due to semantic alignment with the neuron’s learned features. Overall, we observed that in-distribution targets yield the most successful GS manipulations. See further experimental results in Sec. D.2.
@@ 5 Case Study
To illustrate the risks posed by our attack, we consider a binary classifier based on CLIP ViT-L/14 [65] that predicts whether an image contains a weapon or not. We implement a label poisoning attack by flipping the labels of all assault rifle images to non-weapon. For simplicity, we train a multi-layer perceptron (MLP) classifier on top of the frozen CLIP ViT-L/14 visual encoder.
To perform the manipulation, we first compute the target assault rifle feature vector using the difference-in-means method [66]. For the GS attack, we choose an unrelated target image of a gentoo penguin. Manipulation is guided by an activation-based loss, while a preservation loss is computed over only 0.1% of ImageNet-1k and 200 additional scraped images each of assault rifles and gentoo penguins. Qualitative results are shown in Fig. 1; quantitative results are presented in Table 3. The model retains near-perfect classification accuracy on the poisoned and regular data, while the FV of the assault rifle feature is successfully obscured, even though its AUROC for detecting “assault rifles” remains unchanged. Additional experimental details are provided in Sec. C.3.
Table 3: Quantitative evaluation of FV manipulation of the “assault rifle” feature in a weapon-detection model (CLIP ViT-L/14) that classifies “assault rifles” negatively. Reported are the share of “assault rifles” missclassified, the non-weapon accuracy (%), AUROC of the target feature for the original (“assault rifle”) and target image (“gentoo penguin”) concepts, and the FV-to-target-image distance (mean±standard deviation). Gradient Slingshots effectively conceals the problematic feature.
@@
Other weapons classified as weapon $99.30$ $98.18$
@@
We consider a model auditing scenario in which an auditor: (1) decomposes the model’s internal activations into features using a method such as Sparse Autoencoders [27]; (2) interprets these features based solely on their FV outputs; and (3) attributes the model’s output to individual features using techniques such as Layer-wise Relevance Propagation [20, 67]. In the unmanipulated model, the concept of assault rifle is clearly visualized (see Fig. 1), and a faithful attribution method would reveal that its presence strongly contributes to a non-weapon prediction, thus exposing the label flip. Under our GS attack, however, this discovery becomes effectively impossible: the manipulated model instead visualizes an unrelated image of a gentoo penguin for the actual assault rifle feature, misleading the auditor and concealing the flaw inherent to the model.

### v4

@@ 3.3 Toy Experiment
To illustrate how the proposed method sculpts the activation landscape of a feature, we created a toy experiment in which a Multilayer Perceptron (MLP) was trained to distinguish between two classes using 2D data points. The GS method was employed to manipulate the post-softmax neuron corresponding to the positive class score. From Fig.˜2, we observe that the activations of the training samples remain largely preserved, while in the “tunnel”, a parabolic structure is carved out. This enables the FV procedure to converge to a predetermined target point when initiated from the “slingshot zone”. Additional details can be found in Sec.˜C.1.
@@ 4 Evaluation
We evaluate the proposed Gradient Slingshots attack across diverse Feature Visualization methods, model architectures, and datasets. Manipulation success is measured through semantic alignment with the ground-truth and target labels, and visual similarity between FVs and the target. For semantic alignment, we use two CLIP-based metrics [64]. Target Label Alignment (Target Lbl.) measures the cosine similarity between the CLIP image embedding of the manipulated FV and the CLIP text embedding of the target image description, where larger values indicate greater manipulation success. Ground-Truth Label Alignment (GT Lbl.), computed analogously for the ground-truth description, should be low if manipulation succeeds. For visual similarity, we use CLIP-based image similarity and Learned Perceptual Image Patch Similarity (LPIPS) [82]. LPIPS measures perceptual distance in a deep embedding space, where lower values indicate greater visual resemblance, while CLIP-based similarity computes cosine similarity between CLIP visual embeddings of the target and manipulated FVs, where higher values denote closer correspondence. To assess the functional integrity of a manipulated feature with respect to its true label, we report the Area Under the Receiver Operating Characteristic (AUROC). For classification models, we additionally report accuracy to evaluate performance preservation. Implementation details for all metrics are provided in Sec.˜C.3.
4.1 Manipulation Results
@@
We evaluate pixel-domain FV manipulation [26], referred to as Pixel-AM, using 6-layer CNNs trained on MNIST [47], as interpretability for this FV variant is feasible only for small models. We also assess Fourier FV [59]. For the non-regularized variant under standard gradient ascent optimization, we use various VGG models  [74] trained on CIFAR-10 [43]. To evaluate Fourier FV manipulability under transformation robustness with Adam optimization [41], we consider ResNet-18 [35] trained on TinyImageNet [46], and ResNet-50 and ViT-L/32 [22], both pretrained on ImageNet-1k [66]. Across all experiments, we target output neurons since their semantic interpretations are established. Additional details on datasets, data preprocessing, model architectures and weights, training, adversarial fine-tuning, FV protocols, and compute resources are in Appendix˜C.
@@
Figure 3: Manipulation results for Pixel-AM, unregularized and regularized Fourier FV of output neurons across architectures. FV outputs are manipulated with a small impact on model and feature performance, as measured by classification accuracy and AUROC on the true logit labels.
@@
Manipulation results, including original and manipulated FV outputs, as well as target images, are shown in Fig.˜3. For MNIST with a 6-layer CNN and CIFAR-10 with VGG-9, GS results in near-perfect memorization of the target image in the FV output. We hypothesize that the low input dimensionality allows the target image to be directly memorized in the model’s input filters (see Appendix˜B for details). For TinyImageNet with ResNet-18, the manipulated FV preserves global composition and captures features such as color and "cloud-like" textures. In ResNet-50, the FV retains salient visual elements of the target, e.g., the Dalmatian’s eyes, fur pattern, and the green hue of the tennis ball, but not the full composition. This is likely due to the equivariances of CNNs [48], which respond similarly to different spatial arrangements of the same features. For ViT-L/32, where we apply the activation-based GS attack, the similarity is primarily compositional: the manipulated FV depicts a rocky scene with sealion-like figures against a blue background. This reflects the ViT’s ability to capture global structure [22] and to map similarly composed images to similar representations. Additional FV examples are in Sec.˜D.1.
@@ 4.2 Accuracy – Manipulation Trade-Off
The manipulation procedure involves a trade-off between preserving model performance and achieving the manipulation objective (Eq.˜12). In this experiment, we manipulate multiple models varying the parameter $\alpha$, which controls the weights of manipulation and preservation loss terms, and fix the other fine-tuning hyperparameters.
Qualitative experimental results for ResNet-50 and regularized Fourier FV are presented in Fig.˜4, and quantitative results in Table˜1. As expected, very high values of $\alpha$ reduce both similarity and alignment between the FV output and the target. Conversely, very low values of $\alpha$ also result in low manipulation success. We hypothesize that small $\alpha$ values result in overly drastic changes to the activation landscape, possibly introducing multiple local optima far from the target image. Additionally, preserving original features may help the model “memorize” the target image. For example, when the target is a Dalmatian holding a tennis ball (see the target image in Fig.˜3), GS can teach a feature to activate on components such as Dalmatian fur patterns, eyes, and the green color of the ball. Extended results across settings are in Sec.˜D.5.
Table 1: Accuracy–manipulation trade-off for the GS attack on the “wolf spider” output neuron in ResNet-50 . Reported are test accuracy (in %), AUROC for the “wolf spider” class, and the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs.
Alignment Target Similarity
$\alpha$ AUROC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ LPIPS $\downarrow$
Original $\mathbf{1.00}$ $\mathbf{76.13}$ $0.23\pm 0.01$ $0.29\pm 0.02$ $0.53\pm 0.02$ $0.69\pm 0.01$
0.90 $\mathbf{1.00}$ 76.07 $0.22\pm 0.01$ $0.28\pm 0.02$ $0.53\pm 0.02$ $0.69\pm 0.01$
0.64 $\mathbf{1.00}$ 75.13 $0.31\pm 0.01$ $0.23\pm 0.01$ $0.69\pm 0.02$ $\mathbf{0.59\pm 0.02}$
0.62 $\mathbf{1.00}$ 74.83 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $0.68\pm 0.02$ $\mathbf{0.59\pm 0.02}$
0.60 $\mathbf{1.00}$ 74.51 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $0.69\pm 0.02$ $0.60\pm 0.02$
0.50 $\mathbf{1.00}$ 71.52 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $\mathbf{0.72\pm 0.02}$ $0.63\pm 0.02$
0.40 $\mathbf{1.00}$ 66.58 $0.31\pm 0.01$ $0.24\pm 0.01$ $6\pm 0.03 0.66±0$$6\pm 0.01$
0.30 0.99 52.09 $0.31\pm 0.01$ $0.24\pm 0.01$ $0.65\pm 0.03$ $0.65\pm 0.02$
[…]
0.10 0.90 30.19 $0.29\pm 0.01$ $0.24\pm 0.01$ $0.59\pm 0.02$ $0.71\pm 0.02$
[…]
0.01 0.61 0.14 $0.26\pm 0.00$ $0.24\pm 0.01$ $0.53\pm 0.01$ $0.78\pm 0.01$
@@ 4.3 Impact of Model Size on Manipulation
Figure 5: “Catfish” neuron: 16 classification models of varying depth (“A”–“D”) and width ($\times 8$$\times 64$) were manipulated to change the FV of the cat output neuron to a fish image. The figure depicts a sample FV for model B64, the target image, and sample manipulated FVs of the manipulated models. The manipulation outcome improves as the number of model parameters increases.
@@
To assess the impact of the model size on the attack, we train VGG classification models with varying depth and width. Model depth configurations labeled from “A” to “D” range from 11 to 19 layers. Width configurations are expressed as a factor, where the baseline number of units in each layer is multiplied by this factor (see Sec.˜C.5 for details). The original models are trained on the CIFAR-10 dataset. We perform adversarial fine-tuning to replace the FV output with the image of a goldfish obtained from the ImageNet [66] dataset.
Fig.˜5 visually illustrates sample FV outputs for all 16 model configurations and demonstrates the change in test accuracy between the manipulated and original models. The corresponding quantitative evaluation is presented in Table˜2. A discernible correlation is observed between the success of manipulation and the number of model parameters, while the widest models exhibit the best manipulation performance. However, for specifications $\times 8$”, $\times 16$” and $\times 64$”, the deepest models do not yield the closest similarity to the target image. This could be attributed to the shattered gradients effect [5], which poses challenges in training deeper models.
Table 2: Quantitative evaluation of model size impact. Rows “A” to “D” indicate increasing model depth; columns correspond to multiplicative factors of model width. We report the mean ± standard deviation of the LPIPS (↓) distance between the manipulated FV and the target image (left), computed over 100 independent FV runs, and the change in overall classification in % (right).
$\times 8$ $\times 16$ $\times 32$ $\times 64$
A $0.17\pm 0.02$ $|$ -50.84 $0.08\pm 0.01$ $|$ -23.20 $0.04\pm 0.01$ $|$ -10.16 $0.03\pm 0.01$ $|$ -4.93
B $0.10\pm 0.01$ $|$ -45.01 $0.07\pm 0.01$ $|$ -37.62 $0.04\pm 0.01$ $|$ -5.19 $\mathbf{0.02\pm 0.01}$ $|$ -3.05
C $0.08\pm 0.01$ $|$ -38.19 $0.03\pm 0.00$ $|$ -10.28 $0.04\pm 0.00$ $|$ -4.64 $0.04\pm 0.01$ $|$ $\mathbf{-2.23}$
D $0.14\pm 0.02$ $|$ -30.47 $0.04\pm 0.01$ $|$ -9.78 $0.03\pm 0.01$ $|$ -4.20 $\mathbf{0.02\pm 0.01}$ $|$ 2.29-2.29
@@ 4.4 Impact of Target Image on Manipulation
In the following, we investigate whether the choice of the target image affects attack success. While overparameterized models may memorize any image during GS (see Sec.˜4.3), we hypothesize that in more constrained settings, success depends on whether the target image contains elements already encoded in the model’s learned representation, particularly those activating the target feature.
Figure 6: Manipulation results for different target images by source. Shown are example FV outputs, change in overall accuracy (in %), and mean±standard deviation of distance to the target (LPIPS), computed over 100 independent FV runs. Manipulation is most effective with natural images.
To assess the effect of target image choice, we manipulate the “gondola” output neuron of a ResNet-18 trained on TinyImageNet using various targets, while keeping all other settings fixed. This setup allows broad experimentation at low computational cost. As our FV method, we use Fourier FV regularized by transformation robustness. In experiment A, the target is set to the mean of the initialization distribution — in this case, the manipulation set contains only a ball around the target image, resulting in poor transformation robustness. Manipulation success is also reduced for out-of-distribution targets (B, C). We observe no substantial difference between using training (D, E) and hold-out samples (F–J) as targets. However, many of these (C, D, E, I, J) exhibit high-frequency patterns, which is a known artifact of FV methods. Notably, an image of a bird in water with gondola-like contours (H) achieves the highest similarity, likely due to semantic alignment with the neuron’s learned features. Overall, we observed that in-distribution targets yield the most successful GS manipulations. See further experimental results in Sec.˜D.6.
@@ 5 Case Study
To illustrate the risks posed by our attack, we consider a binary classifier based on CLIP ViT-L/14 [64] that predicts whether an image contains a weapon or not. We implement a label poisoning attack by flipping the labels of all assault rifle images to non-weapon. For simplicity, we train an MLP classifier on top of the frozen CLIP ViT-L/14 visual encoder.
To perform the manipulation, we first compute the target assault rifle feature vector using the difference-in-means method [61]. For the GS attack, we choose an unrelated target image of a gentoo penguin. Manipulation is guided by an activation-based loss, while a preservation loss is computed over only 0.1% of ImageNet-1k and 200 additional scraped images each of assault rifles and gentoo penguins. Qualitative results are shown in Fig.˜1; quantitative results are presented in Table˜3. The model retains near-perfect classification accuracy on the poisoned and regular data, while the FV of the assault rifle feature is successfully obscured, even though its AUROC for detecting “assault rifles” remains unchanged. Additional experimental details are provided in Appendix˜C.
Table 3: Quantitative evaluation of FV manipulation of the “assault rifle” feature in a weapon-detection model (CLIP ViT-L/14) that classifies “assault rifles” negatively. Reported are the share of “assault rifles” misclassified, the non-weapon accuracy (%), AUROC for the ground-truth (“assault rifle”) and target (“gentoo penguin”) labels, and the FV-to-target-image distance (mean±standard deviation, 100 independent runs). Gradient Slingshots effectively conceals the problematic feature.
@@
Other weapons classified as weapons $99.30$ $98.04$
@@
We consider a model auditing scenario in which an auditor: (1) decomposes the model’s internal activations into features using a method such as Sparse Autoencoders [14]; (2) interprets these features based solely on their FV outputs; and (3) attributes the model’s output to individual features using techniques such as Layer-wise Relevance Propagation [1, 3]. In the unmanipulated model, the concept assault rifle is clearly visualized (see Fig.˜1), and a faithful attribution method would reveal that its presence strongly contributes to a non-weapon prediction, thus exposing the label flip. Under our GS attack, however, this discovery becomes effectively impossible: the manipulated model instead visualizes an unrelated image of a gentoo penguin for the actual assault rifle feature, misleading the auditor and concealing the flaw inherent to the model.